### v1

Abstract
The set of relevant cuts in a graph is the union of all minimum weight bases of the cut space. A cut is relevant if and only if it is the a minimum weight cut between two distinct vertices.
keywords:
[…]
A Short Note on Relevant Cuts
[…]
Nico DomschkeBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanynico@bioinf.uni-leipzig.de
[…]
Thomas GatterBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanythomas@bioinf.uni-leipzig.de
[…]
Peter F. StadlerBioinformatics Group, Department of Computer Science & Interdisciplinary Center for Bioinformatics & Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig & School of Embedded Composite Artificial Intelligence, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germany; Max Planck Institute for Mathematics in the Sciences, Inselstraße 22, D-04103 Leipzig, Germany; Department of Theoretical Chemistry, University of Vienna, Währingerstraße 17, A-1090 Wien, Austria; Facultad de Ciencias, Universidad National de Colombia; Bogotá, Colombia; Santa Fe Institute, 1399 Hyde Park Rd., Santa Fe NM 87501, USAstudla@bioinf.uni-leipzig.de
[…]
05C21,05C50
1 Introduction
Cuts in graphs are closely related to cycles. The cycle space and the cut space form orthogonal subspaces of the $|E|$-dimensional vector space over GF(2) spanned the edges of $G$. Nevertheless, graphs can contain edge sets called bi-cycles that are both cycles and cuts [21]. The cycle space, and in particular bases of the cycle space comprising short cycles, play an important role in cheminformatics [6, 13, 15] and structural biology [14]. In particular, minimum cycle bases composed of cycles with minimal total length or edge weight, and their union, known as the set of relevant cycles [24], are of use [1].
Like cycle bases, cut bases can also be defined in directed graphs, where they are usually referred to as co-cycle bases [12]. In this case, co-cycles are considered as vectors over $\mathbb{Q}$. Co-cycle bases have found applications in particular in crystallography [4, 25].
Minimum cut bases in undirected graphs are studied in detail in [3], their counterparts in directed graphs were analyzed in [11, 12], highlighting the special role of Gomory-Hu trees. These are (not necessarily spanning) trees on $V(G)$ with edge weights $\lambda_{e}$ such that for each pair $s,t\in V(G)$, the minimum weight $\lambda_{e}$ along the path between $s$ and $t$ in $T$ equals the weight of a minimum $s,t$-cut [7]. Since Gomory-Hu trees exist for all graphs, certain sets of $n-1$ distinct minimal cuts already provide a minimal cut for any pair of distinct vertices. As shown in [3, 11], the cuts defined by a Gomory-Hu tree form minimal cut (co-cycle) bases. Moreover, they are weakly fundamental, i.e., their cuts can be ordered in such a way that every cut contains an edge that is not present in any of its predecessors.
The construction principle of Gomory-Hu trees ensures that the associated cuts are nested, i.e., for any two distinct cuts $(A,\bar{A})$ and $(B,\bar{B})$, one of the intersections $A\cap B$, $A\cap\bar{B}$, $\bar{A}\cap B$ and $\bar{A}\cap\bar{B}$ is empty. A simplified algorithm [8], however, alleviates this restriction, suggesting that minimal cut bases are more general than the Gomory-Hu bases.
2 Preliminaries
Throughout, $G=(V,E,w)$ is an undirected connected graph with edge weights $w:E\to\mathbb{R}$. In the context of chemical chemical application, integer weight corresponding to the order of a chemical bond will be most relevant. A cut is a proper bipartition $(V^{\prime},V^{\prime\prime})$ of the vertex set $V(G)$, i.e., $V^{\prime}\neq\emptyset$, $V^{\prime\prime}\neq\emptyset$, $V^{\prime}\cap V^{\prime\prime}=\emptyset$, and $V^{\prime}\cup V^{\prime\prime}=V$. A cutset $D\coloneqq\operatorname{cut}(V^{\prime},V^{\prime\prime})=\{xy\in E|x\in V^{\prime},y\in V^{\prime\prime}\}$. The weight of a cutset is $|D|\coloneqq\sum_{xy\in D}w(xy)$. In the following we will usually write $\partial V^{\prime}=\partial V^{\prime\prime}=\operatorname{cut}(V^{\prime},V^{\prime\prime})=\operatorname{cut}(V^{\prime\prime},V^{\prime})$. In particular, the edges incident to each vertex $x\in V$ form a cutset $\partial\{x\}$. A cutset $D\in\mathfrak{D}$ that does not include another cutset $D^{\prime}$ of $G$ as a proper subset is called a bond of $G$. Clearly, each $\partial\{x\}$ is bond. Let $T$ be a spanning tree of $G$, and $e$ an edge of $T$. Removal of $e$ decomposed $T$ into exactly two connected components with vertex sets $V_{e}^{\prime}$ and $V_{e}^{\prime\prime}$. By construction the induced subgraphs $G[V_{e}^{\prime}]$ and $G[V_{e}^{\prime\prime}]$ are connected. The cut $\partial V_{e}^{\prime}$ is also a bond.
The cutsets of $G$ form a vector space $\mathfrak{D}\coloneqq\mathfrak{D}(G)$ over GF(2) called the cut space. We write $\oplus$ for the vector addition and note that this corresponds to the symmetric difference of edge sets. Thus we have $D\oplus D=\varnothing$. It is not difficult to check, moreover, that
$\partial W=\bigoplus_{x\in W}\partial\{x\}$ (1)
For connected graphs, $\mathfrak{D}$ has dimension $n-1$, where $n=|V|$ is the number of vertices in $G$. In particular, each of the sets $\mathcal{D}_{x_{0}}\coloneqq\{\partial\{x\}|x\in V\setminus\{x_{0}\}\}$ for any $x_{0}\in V$, and each of the sets $\mathcal{D}_{T}\coloneqq\{\partial V_{e}^{\prime}|e\in E(T)|\}$ for any spanning tree $T$ of $G$ is basis of the cut space, see e.g. [2].
A cut $(V^{\prime},\bar{V}^{\prime})$ is an $s,t$-cut if $s\in V^{\prime}$ and $t\in\bar{V}^{\prime}$. Clearly, the cut space $\mathfrak{D}$ contains an $s,t$-cut for every pair $s\neq t$ of distinct vertices. Below, we will make use of the fact that this is also true for every basis of $\mathfrak{D}$, see e.g. also [3]. For completeness we include a short proof.
Lemma 2.1.
Let $G$ be a connected graph, $s,t\in V(G)$ distinct vertices, and $\mathcal{D}$ a cut basis. Then $\mathcal{D}$ contains an $s,t$-cut.
Proof 2.2.
Let $\mathcal{D}=\{\partial U_{i}|i=1,\dots,n-1\}$ be a cut-basis, and set $Z\coloneqq\bigcup_{i}U_{i}$. Then any cut that can be obtained as $\oplus$-sum of a subset of $\partial U_{i}$ does not contain any edge with both endpoints in $Z$ and thus must be of the form $(W^{\prime},\bar{W}^{\prime})$ with $Z\subseteq W^{\prime}$ or $Z\subseteq\bar{W}^{\prime}$. Assume $s,t\in Z$. Then the cut space generated by $\mathcal{D}$ does not contain an $s,t$-cut. This is impossible, however, since $\partial\{s\}$ is an $s,t$-cut is contained in the cut space, contradicting the assumption that $\mathcal{D}$ is a basis.
Definition 2.3.
A basis $\mathcal{D}$ of the cut space $\mathfrak{D}$ is a minimum cut basis if $\sum_{D\in\mathcal{D}}|D|$ is minimum.
Definition 2.4.
Let $G=(V,E,w)$ be a weighted graph and let $T^{*}=(V,F,\lambda)$ be an edge-weighted tree on the same vertex set. Then $T^{*}$ is a Gomory-Hu tree for $G=(V,E,w)$ if
  • (G1)
    for every $st\coloneqq e\in F$, the bipartition of $V$ induced by removing $e$ from $F$ defines a cut of $G$, which we denote by $(V_{e},\bar{V}_{e})$;
  • (G2)
    for every $st\coloneqq e\in F$, the weight $\lambda_{e}$ of $e$ is the weight of a minimum $s,t$-cut;
  • (G3)
    for all $u,v\in V$, the weight of a minimum $u,v$-cut in $G$ equals the minimum weight $\lambda_{e}$ along the unique path $P_{uv}$ connecting $u$ and $v$ in $T^{*}$.
A key result in the theory of flow networks is
Proposition 2.5.
[7] Every weighted graph $G=(V,E,w)$ has a Gomory Hu tree $T^{*}=(V,F,\lambda^{*})$.
Proposition 2.6.
[3] The cut-sets $\partial{V_{e}}$ induced by the edges $e\in F$ of a Gomory-Hu tree $T^{*}$ for $G$ form a minimum cut basis of $G$.
3 Relevant Cuts
The notion of relevant cycles was and studied for cycles bases of undirected graphs in [24]. Interestingly, the notion was proposed already in 1971 by Morris Plotkin in the context of chemical ring perception [17]. More generally, the concept of relevant elements pertains to matroids with a weight function in general and reads as follows for the special case of cut spaces:
Definition 3.1.
A cut $D\in\mathfrak{D}$ is relevant if there is a minimum cut basis $\mathcal{D}$ of $\mathfrak{D}$ such that $D\in\mathcal{D}$.
The following simple characterization of relevant elements holds for matroids in general [24]:
Lemma 3.2.
A cut $D\in\mathfrak{D}$ is relevant if and only if it cannot be written as an $\oplus$ sum of shorter cuts $D_{1},\dots,D_{k}\in\mathfrak{D}$, i.e., such that $|D_{i}|<|D|$ for $1\leq i\leq k$.
Proof 3.3.
Since $\mathfrak{D}$ is a vector space it is also matroid. Suppose $D$ is contained in the minimum cut basis $\mathcal{D}$. If $D$ can be written as the $\oplus$-sum of shorter cuts $D_{1}$ through $D_{k}$, it could be exchanged with one of these cuts, contradicting minimality of $\mathcal{D}$. Now suppose that $D$ cannot be written as a sum shorter cuts. Then $D=D_{1}\oplus\dots\oplus D_{k}$ and there is at least one cut $D_{i}$ with $|D_{i}|\geq|D|$. Now $D_{i}=D_{1}\oplus\dots\oplus D_{i-1}\oplus D\oplus D_{i+1}\oplus\dots\oplus D_{k}$ and thus $\mathcal{D}^{\prime}=\{D\}\cup\mathcal{D}\setminus\{D_{i}\}$ is again a cut basis. By minimality of $\mathcal{D}$ we must have $|D_{i}|=|D|$ and thus $\mathcal{D}^{\prime}$ is also a minimum cut basis. It follows that $D$ is relevant.
In particular, therefore, all weight-minimal cuts are relevant, see also [20]. On the other hand, we can rule out cuts that lead to more than two connected subgraphs:
Lemma 3.4.
If $D\in\mathfrak{D}$ is relevant, then $D$ is a bond.
Proof 3.5.
Suppose, for contradiction, that $D$ is not a bond. Then there is a cutset $D^{\prime}\in\mathfrak{D}$ such that $D^{\prime}\subsetneq D$. Moreover, since the cutsets form a vector space, $D^{\prime\prime}\coloneqq D\oplus D^{\prime}=D\setminus D^{\prime}\in\mathfrak{D}$. By construction we have $D=D^{\prime}\oplus D^{\prime\prime}$ and $|D^{\prime}|<|D|$, and $|D^{\prime\prime}|<|D|$, and thus $D$ is not a relevant cut.
The cuts induced by the edges of a Gomory Hu tree are nested and form a basis of the cut space. The cut-set $\partial A$ of any cut $(A,\bar{A})$ of $G=(V,E,w)$ therefore can be expressed as a linear combination of the cut sets defined by a Gomory-Hu tree $T^{*}$. More precisely,
$\partial A=\bigoplus_{\begin{subarray}{c}uv=e\in F\\ u\in A,v\in\bar{A}\end{subarray}}\partial V_{e}$ (2)
where $\partial V_{e}$ denotes the cutset of the cut $(V_{e},\bar{V}_{e})$ obtained by removing $e$ from $T^{*}$. In other words, $\partial A$ is the $\oplus$-sum of the cuts defined by edges in the Gomory-Hu tree $T^{*}$ with one vertex in $A$ and the other one in $\bar{A}$. Thus $u$ and $v$ are both in $A$ or both in $\bar{A}$ if and only if they are separated by an even number of cut edges along the unique path $P_{uv}$ in $T^{*}$.
Theorem 3.6.
A cut $\partial U$ in $G$ is relevant if and only if it is a weight-minimal $u,v$-cut for some pair of distinct vertices $u$ and $v$.
Proof 3.7.
Claim: Every weight-minimal $u,v$-cut is relevant.
Let $D^{*}_{uv}$ be an arbitrary but fixed weight-minimal $u,v$-cut. Sort the list of all cuts in $G$ by non-decreasing weight such that $D^{*}_{uv}$ is the first in the list with its weight. Thus no shorter cut is an $u,v$-cut. Running the canonical greedy algorithm until the list entry before $D^{*}_{uv}$ produces a set $\mathcal{B}^{\prime}$ of linearly independent cuts $\partial U_{i}$ such that $u,v\in U_{i}$ and no $u,v$-cut can be generated. Therefore $\mathcal{B}^{\prime}\cup\{D^{*}_{uv}\}$ is again linearly independent, i.e., the canonical greedy algorithm includes $D^{*}_{uv}$ in the growing basis and hence $D^{*}_{uv}$ is relevant. $\triangleleft$
Claim: If $A$ is not a minimum weight cut for any pair of vertices $u$ and $v$, then it is not relevant.
Let $T^{*}$ be an arbitrary Gomory-Hu tree for $G$. According to equ.(2), $A$ is a linear combination of cut sets $\partial V_{e}$ for an arbitrary Gomory Hu tree $T^{*}$ for the edges that separate vertice $u\in V_{e}$ and $v\in\bar{V}_{e}$. By (G2) and (G3), the weight $\lambda_{e}$ of the edge $e=uv$ in $T^{*}$ satisfies $\lambda_{e}=|\partial V_{e}|$ and, by construction, is the minimum weight of cuts separating $u$ and $v$. Since $A$ is not a minimum weight cut for any pair of vertices, it is in particular not a minimum weight cut for $u$ and $v$. Thus we have $|\partial V_{e}|<|A|$ for every edge $e$ of the $T^{*}$, i.e., $D$ is a linear combination of strictly shorter cuts. By Lemma 3.2, therefore, $A$ it is not relevant. $\triangleleft$
Taken together a cut $D\in\mathfrak{D}$ is relevant if and only if its is a minimum weight cut for at least one pair of vertices $u,v\in V$.
There are graphs with an exponential number of minimal $s,t$-cuts. For instance the $K_{2,n-2}$ with unit edge-weights and $\{s,t\}$ for one part of the bipartition [16]. Since all these cuts are relevant by Thm. 3.6, the number of relevant cuts, like the number of relevant cycles [24], may also grow exponentially with the size of the graph $G$. In the unweighted case, i.e., $w_{pq}=1$ for all edges $pq\in E$, the number of relevant cut is bounded by $n^{\Delta}$, where $\Delta$ is the maximum degree because every minimal $s,t$-cut contains at most $\Delta$ edges and there are no more than $O(n^{\Delta})$ distinct edge sets with at most $\Delta$ edges.
This argument does not generalize to arbitrary weights, however: Consider the graph $G$ obtained from two copies of binary trees with $n$ leaves by identifying corresponding leaves. Thus $G$ has $23n-$ 2 vertices and a maximum vertex degree of $3$. Suppose all edge weights are unity except for edges incident with the leaves, which have weight $1/(2n)$. Denoting the roots of the trees by $s$ and $t$, every cut that contains an edge not incident with a tree-leaf has weight $>1$, while cutting exactly one of the two edges incident with a leaf yields an $s,t$-cut with weight $1/2$. There are $2^{n}$ distinct cuts of this type, all of which a minimum weight $s,t$-cuts.
4 Computational Considerations
By Thm. 3.6, the most direct approach to computing the set of relevant cuts is to compute, for all distinct pairs $s,t\in V$, the set of all minimal $s,t$-cuts. Even though there may be exponentially many relevant cuts, it is possible to encode them in an intermediary data structure that can be constructed in polynomial time. As shown in [9], the set all minimal $a,b$-cuts for all $a,b$, i.e., the set of relevant cycles, can be computed from a Gomory-Hu tree $T^{*}$ and the Picard-Queyranne (PQ) DAGs for the edges $st$ of $T^{*}$. The PQ DAG $D(s,t)$ is a compact representation of all minimal $s,t$-cuts [16], containing the pertinent cuts as the closed sets of the DAG. Both $T^{*}$ and the PQ-DAG $D(st)$ for any $st$ can be computed in polynomial time. From this combined data structure, $\mathcal{K}$ can be listed with $O(|\mathcal{K}|n)$ effort [9]. This is reminiscent of the polynomially many “prototypes” for relevant cycles described in [24], from which all relevant cycles can be listed. There is an important difference, however: For relevant cycles, there is a polynomial-time algorithm to count their number based on counting a constrained set of shortest paths [24]. In contrast, counting the number of minimal $s,t$-cuts is #P-complete [18], and thus at least as hard as NP-complete decision problems [22].
In addition to the approach of [9], several other algorithms for listing $\mathcal{K}$ are feasible. By Lemma 3.4, one could use the bonds of $G$ as a candidate set and extract $\mathcal{K}$ using the canonical greedy algorithm similar to Horton’s approach to computing minimum cycle bases [10]. Very recently, an algorithm for listing all $(s,t)$-bonds with near linear delay [19] has become available. A disadvantage is that the set of bonds must first be generated and then sorted by non-decreasing weight.
A more appealing alternative is to use one of the approaches for ordered generation of cuts. The first method of this type was proposed in [23]. Later an improved version based on a different solutions of maximum flow problems was published [26]. Since these algorithm produce cuts in the order of non-increasing weights, it is not difficulty to whether a cut $(A,\bar{A})$ is a minimum weight cut for any pair $u\in A$ and $v\in A^{\prime}$. This can be done by keeping a matrix $W$, initialized by $W_{st}=\infty$ for all $s,t$, and updated by the following rule: if $W_{u,v}=\infty$, $u\in A$ and $v\in\bar{A}$ then $W_{u,v}\leftarrow|\partial A|$. In particular, it is not necessary to check for linear independence of a cut with all cuts of strictly smaller weight as is the case for minimal cycle bases. Moreover, the generation process can be terminated as soon as (i) minimum weight cuts have been observed for all vertices, i.e., upon generation of $(A,\bar{A})$, no entries $\infty$ were present in $W$, and (ii) the weight of the current cut is strictly larger than the weight of the previous one, which coincides with the largest finite entry in $W$. The enumeration algorithm in [26] has $\tilde{O}(nm)$ delay, which dominates the $O(n^{2})$ effort for “filtering out” cuts that are not minimal $u,v$-cuts for any pair $u,v$.
5 Concluding Remarks
Beyond the interest in relevant cuts implied by the analogy to relevant cycles, they are also of practical interest e.g. for the design of well-behaved crossover operators in graph-based genetic algorithms. Recently, cut-and-join crossover operators for chemical graphs have been investigated and benchmarked in detail for search tasks in “chemical space”, demonstrating their efficacy [5]. It remains an open question, however, whether the restriction of cuts the set relevant cuts remains sufficient to ensure global reachability in the space of chemical graphs in conjunction with degree-preserving or even more restricted join operations.
Acknowledgements
This work was supported by the German Research Foundation (DFG) within SPP2363 (460865652) and the Novo Nordisk Foundation as part of MATOMIC (0066551). Research in the Stadler lab is supported by the German Federal Ministry of Education and Research BMBF through DAAD project 57616814 (SECAI, School of Embedded Composite AI).