### v3

Abstract
The set of relevant cuts in a graph is the union of all minimum weight bases of the cut space. We show that a cut is relevant if and only if it is a minimum weight cut between two distinct vertices. Moreover, we give a characterization in terms of Picard-Queyranne Directed Acyclic Graphs that can be used to accelerate the enumeration of the relevant cuts. Finally, we perform an experimental evaluation by comparing with state-of-the-art algorithms.
keywords:
Cut Space; Matroid; Gomory-Hu tree; Enumeration;
[…]
A Short Note on Relevant Cuts
[…]
Nico DomschkeBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanydnico@bioinf.uni-leipzig.de
[…]
Thomas GatterBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanythomas@bioinf.uni-leipzig.de
[…]
Richard GolnikBioinformatics Group, Department of Computer Science, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germanyrichard@bioinf.uni-leipzig.de
[…]
The Students of The Advanced Bioinformatics Course 2024Department of Computer Science, Leipzig University, Augustusplatz 10, D-04109 Leipzig, Germany
[…]
Peter F. StadlerBioinformatics Group, Department of Computer Science & Interdisciplinary Center for Bioinformatics & Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig & School of Embedded Composite Artificial Intelligence, Leipzig University, Härtelstraße 16–18, D-04107 Leipzig, Germany; Max Planck Institute for Mathematics in the Sciences, Inselstraße 22, D-04103 Leipzig, Germany; Department of Theoretical Chemistry, University of Vienna, Währingerstraße 17, A-1090 Wien, Austria; Facultad de Ciencias, Universidad National de Colombia; Bogotá, Colombia; Santa Fe Institute, 1399 Hyde Park Rd., Santa Fe NM 87501, USAstudla@bioinf.uni-leipzig.de
[…]
05C21,05C50
1 Introduction
Cuts in graphs are closely related to cycles. The cycle space and the cut space form orthogonal subspaces of the $|E|$-dimensional vector space over GF(2) spanning the edges of a graph $G=(V,E)$. Nevertheless, graphs can contain edge sets called bi-cycles that are both cycles and cuts [27]. The cycle space, and in particular bases of the cycle space comprising short cycles, play an important role in cheminformatics [8, 17, 20] and structural biology [19, 18]. In particular, minimum cycle bases composed of cycles with minimal total length or edge weight, and their union, known as the set of relevant cycles [30], are of use [3].
Like cycle bases, cut bases can also be defined in directed graphs, where they are usually referred to as co-cycle bases [16]. In this case, co-cycles are considered as vectors over $\mathbb{Q}$. Co-cycle bases have found applications in particular in crystallography [6, 31].
Minimum cut bases in undirected graphs are studied in detail in [5], their counterparts in directed graphs were analyzed in [15, 16], highlighting the special role of Gomory-Hu trees. These are (not necessarily spanning) trees on $V(G)$ with edge weights $\lambda_{e}$ such that for each pair $s,t\in V(G)$, the minimum weight $\lambda_{e}$ along the path between $s$ and $t$ in $T$ equals the weight of a minimum $s,t$-cut [11]. Since Gomory-Hu trees exist for all graphs, certain sets of n1n-1 distinct minimal cuts already provide a minimal cut for any pair of distinct vertices. As shown in [5, 15], the cuts defined by a Gomory-Hu tree form minimal cut (co-cycle) bases. Moreover, they are weakly fundamental, i.e., their cuts can be ordered in such a way that every cut contains an edge that is not present in any of its predecessors.
The construction principle of Gomory-Hu trees ensures that the associated cuts are nested, i.e., for any two distinct cuts $(A,\bar{A})$ and $(B,\bar{B})$, one of the intersections $A\cap B$, $A\cap\bar{B}$, $\bar{A}\cap B$ and $\bar{A}\cap\bar{B}$ is empty. A simplified algorithm [13], however, alleviates this restriction, suggesting that minimal cut bases are more general than the Gomory-Hu bases.
Beyond the interest in relevant cuts implied by the analogy to relevant cycles, these are also of practical interest, e.g., for the design of well-behaved crossover operators in graph-based genetic algorithms. For example, recently, cut-and-join crossover operators for chemical graphs have been investigated and benchmarked in detail for search tasks in “chemical space”, demonstrating their efficacy [7].
In the following sections we show that relevant cuts coincide with the cuts that are weight-minimal for some pair of vertices; these results follows from previously known properties of minimal cut bases (see Sections 2 and 3). Then, we use these insights to construct novel algorithms for computing a cut basis and the set of relevant cuts of arbitrary graphs (Section 4). Moreover, we benchmark pre-existing and novel algorithms on chemical and random graphs (Section 5). Finally, we draw our conclusions (Section 6).
2 Preliminaries
Throughout, $G=(V,E,w)$ is an undirected connected graph with edge weights $w:E\to\mathbb{R}$. In the context of chemical applications, integer weights corresponding to the order of a chemical bond will be most relevant. A cut is a proper bipartition $(V^{\prime},V^{\prime\prime})$ of the vertex set $V(G)$, i.e., $^{\prime}\neq\emptyset, V′′V^{\prime\prime}\neq\emptyset$, $V^{\prime}\cap V^{\prime\prime}=\emptyset$, and $V^{\prime}\cup V^{\prime\prime}=V$. A cutset $D\coloneqq\operatorname{cut}(V^{\prime},V^{\prime\prime})=\{xy\in E|x\in V^{\prime},y\in V^{\prime\prime}\}$. The weight of a cutset is $|D|\coloneqq\sum_{xy\in D}w(xy)$. In the following we will usually write $\partial V^{\prime}=\partial V^{\prime\prime}=\operatorname{cut}(V^{\prime},V^{\prime\prime})=\operatorname{cut}(V^{\prime\prime},V^{\prime})$. In particular, the edges incident to each vertex $x\in V$ form a cutset $\partial\{x\}$. A cutset $D$ that does not include another cutset $D^{\prime}$ of $G$ as a proper subset is called a bond of $G$. Clearly, each $\partial\{x\}$ is a bond. Let $T$ be a spanning tree of $G$, and $e$ an edge of $T$. Removal of $e$ decomposes $T$ into exactly two connected components with vertex sets $V_{e}^{\prime}$ and $V_{e}^{\prime\prime}$. By construction the induced subgraphs $G[V_{e}^{\prime}]$ and $G[V_{e}^{\prime\prime}]$ are connected. The cut $\partial V_{e}^{\prime}$ is also a bond.
The cutsets of $G$ form a vector space $\mathfrak{D}\coloneqq\mathfrak{D}(G)$ over GF(2) called the cut space. We write $\oplus$ for the vector addition and note that this corresponds to the symmetric difference of edge sets. Thus we have $D\oplus D=\varnothing$. It is not difficult to check, moreover, that
$\partial W=\bigoplus_{x\in W}\partial\{x\}$ (1)
for any $W\subseteq V$. For connected graphs, $\mathfrak{D}$ has dimension n1n-1, where $n=|V|$ is the number of vertices in $G$. In particular, each of the sets $\mathcal{D}_{x_{0}}\coloneqq\{\partial\{x\}|x\in V\setminus\{x_{0}\}\}$ for any $x_{0}\in V$, and each of the sets $\mathcal{D}_{T}\coloneqq\{\partial V_{e}^{\prime}|e\in E(T)\}$ for any spanning tree $T$ of $G$ is basis of the cut space, see e.g. [4].
A cut $(V^{\prime},\bar{V}^{\prime})$ is an $s,t$-cut if $s\in V^{\prime}$ and $t\in\bar{V}^{\prime}$. Clearly, the cut space $\mathfrak{D}$ contains an $s,t$-cut for every pair $s\neq t$ of distinct vertices. Less obviously, this is also true for every basis $\mathcal{D}$ of $\mathfrak{D}$, see, e.g., [5]. For completeness, we include a short proof of this statement, which will be used repeatedly throughout this contribution:
Lemma 2.1.
Let $G$ be a connected graph, $s,t\in V(G)$ distinct vertices, and $\mathcal{D}$ a cut basis. Then $\mathcal{D}$ contains an $s,t$-cut.
Proof 2.2.
Let $\mathcal{D}=\{\partial U_{i}|i=1,\dots,n-1\}$ be a cut-basis comprising cuts $(U_{i},V\setminus U_{i})$ and assume that $\mathcal{D}$ does not contain an $s,t$-cut. Without loosing generality, we may therefore assume that $s,t\in U_{i}$ for all $i$. Consider the set $Z\coloneqq\bigcap_{i}U_{i}$. We have $s,t\in Z$, and moreover, any cut that can be obtained as an $\oplus$-sum of cuts in $\mathcal{D}$ if of the form $(V^{\prime},\bar{V^{\prime}})$ with $Z\subseteq V^{\prime}$ or $Z\subseteq\bar{V^{\prime}}$ since none of the cuts $\partial U_{i}$ contains an edge with both endpoints in $Z$. Hence, note of the cuts $(V^{\prime},\bar{V}^{\prime})$ is an $s,t$-cut, i.e., the $\oplus$-operation on $\mathcal{D}$ cannot generate any $s,t$-cut. The cut space, however, contains in particular the $s,t$-cut $\partial\{s\}$, contradicting our assumption that the $\mathcal{D}$ is a cut basis.
Definition 2.3.
A basis $\mathcal{D}$ of the cut space $\mathfrak{D}$ is a minimum cut basis if $\sum_{D\in\mathcal{D}}|D|$ is minimum.
Definition 2.4.
Let $G=(V,E,w)$ be a weighted graph and let $T^{*}=(V,F,\lambda)$ be an edge-weighted tree on the same vertex set. Then $T^{*}$ is a Gomory-Hu tree for $G=(V,E,w)$ if
  • (G1)
    for every $uv\coloneqq e\in F$, the bipartition of $V$ induced by removing $e$ from $F$ defines a cut of $G$, which we denote by $(V_{e},\bar{V}_{e})$;
  • (G2)
    for every $uv\coloneqq e\in F$, the weight $\lambda_{e}$ of $e$ is the weight of a minimum $s,t$-cut;
  • (G3)
    for all $s,t\in V$, the weight of a minimum $s,t$-cut in $G$ equals the minimum weight $\lambda_{e}$ along the unique path $P_{st}$ connecting $s$ and $t$ in $T^{*}$.
A key result in the theory of flow networks is
Proposition 2.5.
[11] Every weighted graph $G=(V,E,w)$ has a Gomory Hu tree $T^{*}=(V,F,\lambda)$.
Proposition 2.6.
[5] The cutsets $\partial{V_{e}}$ induced by the edges $e\in F$ of a Gomory-Hu tree $T^{*}$ for $G$ form a minimum cut basis of $G$.
While there may be exponentially many minimal $s,t$-cuts for any such fixed node pair, Picard and Queyranne [22] introduced a directed acyclic graph (as a shorthand, we write a PQ-DAG $Q_{s,t}$) that acts as a compact representation and may nevertheless be derived in polynomial time. It is obtained from $G$ by contracting each connected component of the residual flow network to a single node for an arbitrary maximum $s,t$-flow. Accordingly, by construction, vertices in $Q_{s,t}$ represent equivalence classes of vertices w.r.t. the strong connected relationship in the residual flow network, hence sets of nodes from $G$. The “closed” node sets in the DAG $Q_{s,t}$, i.e., those without out-edges, correspond to the minimal $s,t$-cuts.
As shown in [12], given PQ-DAG representations $Q_{u,v}$, corresponding to all minimal $u,v$-cuts, for the edges of the Gomory-Hu tree $uv\in T^{*}$, any PQ-DAG $Q_{s,t}$, where $st$ is not an edge in $T^{*}$, can be computed as a recombination of PQ-DAGs corresponding to edges along the unique path $P_{st}$ connecting $s$ and $t$ in $T^{*}$, improving algorithmic complexity compared to direct construction.
3 Relevant Cuts
The notion of relevant cycles was studied for cycles bases of undirected graphs in [30]. Interestingly, the notion was proposed already in 1971 by Morris Plotkin in the context of chemical ring perception [23]. More generally, the concept of relevant elements pertains to matroids with a weight function in general and reads as follows for the special case of cut spaces:
Definition 3.1.
A cut $D\in\mathfrak{D}$ is relevant if there is a minimum cut basis $\mathcal{D}$ of $\mathfrak{D}$ such that $D\in\mathcal{D}$.
The following simple characterization of relevant elements holds for matroids in general [30]:
Lemma 3.2.
A cut $D\in\mathfrak{D}$ is relevant if and only if it cannot be written as an $\oplus$ sum of shorter cuts $D_{1},\dots,D_{k}\in\mathfrak{D}$, i.e., such that $|D_{i}|<|D|$ for $1\leq i\leq k$.
Proof 3.3.
Since $\mathfrak{D}$ is a vector space it is also matroid. Suppose $D$ is contained in the minimum cut basis $\mathcal{D}$. If $D$ can be written as the $\oplus$-sum of shorter cuts $D_{1}$ through $D_{k}$, it could be exchanged with one of these cuts, contradicting minimality of $\mathcal{D}$. This is justified as follows: We know that at least one cut $D_{1}$ through $D_{k}$ is not element of $\mathfrak{D}$, otherwise $D\notin\mathfrak{D}$ by linear independence. If exactly one cut $D_{i}$ is not an element of $\mathfrak{D}$, $D_{i}$ and $D$ can be exchanged. If multiple cuts are not in $\mathfrak{D}$, we know they can be constructed as $\oplus$-sums of yet shorter cuts, and we apply this argument recursively.
Now suppose that $D$ cannot be written as a sum of shorter cuts. Then $D=D_{1}\oplus\dots\oplus D_{k}$ and there is at least one cut $D_{i}$ with $|D_{i}|\geq|D|$. Now $D_{i}=D_{1}\oplus\dots\oplus D_{i-1}\oplus D\oplus D_{i+1}\oplus\dots\oplus D_{k}$ and thus $\mathcal{D}^{\prime}=\{D\}\cup\mathcal{D}\setminus\{D_{i}\}$ is again a cut basis. By minimality of $\mathcal{D}$ we must have $|D_{i}|=|D|$ and thus $\mathcal{D}^{\prime}$ is also a minimum cut basis. It follows that $D$ is relevant.
In particular, therefore, all weight-minimal cuts are relevant, see also [26]. On the other hand, we can rule out cuts that lead to more than two connected subgraphs:
Lemma 3.4.
If $D\in\mathfrak{D}$ is relevant, then $D$ is a bond.
Proof 3.5.
Suppose, for contradiction, that $D$ is not a bond. Then there is a cutset $D^{\prime}\in\mathfrak{D}$ such that $D^{\prime}\subsetneq D$. Moreover, since the cutsets form a vector space, $D^{\prime\prime}\coloneqq D\oplus D^{\prime}=D\setminus D^{\prime}\in\mathfrak{D}$. By construction we have $D=D^{\prime}\oplus D^{\prime\prime}$ and $|D^{\prime}|<|D|$, and $|D^{\prime\prime}|<|D|$, and thus $D$ is not a relevant cut.
The cuts induced by the edges of a Gomory Hu tree are nested and form a basis of the cut space. The cutset $\partial V^{\prime}$ of any cut $(V^{\prime},\bar{V^{\prime}})$ of $G=(V,E,w)$ therefore can be expressed as a linear combination of the cut sets defined by a Gomory-Hu tree $T^{*}$. More precisely,
$\partial V^{\prime}=\bigoplus_{\begin{subarray}{c}uv=e\in F\\ u\in V^{\prime},v\in\bar{V^{\prime}}\end{subarray}}\partial V_{e}$ (2)
where $\partial V_{e}$ denotes the cutset of the cut $(V_{e},\bar{V}_{e})$ obtained by removing $e$ from $T^{*}$. In other words, $\partial V^{\prime}$ is the $\oplus$-sum of the cuts defined by edges in the Gomory-Hu tree $T^{*}$ with one vertex in $V^{\prime}$ and the other one in $\bar{V^{\prime}}$. Thus $s$ and $t$ are both in $V^{\prime}$ or both in $\bar{V^{\prime}}$ if and only if they are separated by an even number of cut edges along the unique path $P_{st}$ in $T^{*}$.
Theorem 3.6.
A cut $\partial V^{\prime}$ in $G$ is relevant if and only if it is a weight-minimal $s,t$-cut for some pair of distinct vertices $s$ and $t$.
Proof 3.7.
Claim: Every weight-minimal $s,t$-cut is relevant.
Let $D^{*}_{st}$ be an arbitrary but fixed weight-minimal $s,t$-cut. Sort the list of all cuts in $G$ by non-decreasing weight such that $D^{*}_{st}$ is the first in the list with its weight. Thus, no shorter cut is an $s,t$-cut. Using the canonical greedy algorithm [9], a minimal cut basis $\mathcal{B}$ is obtained by iterating over the sorted lists of cuts, adding a cut $D$ to the initial empty set $\mathcal{B}$ if and only if $\mathcal{B}\cup\{D\}$ is linearly independent. Now consider the set $\mathcal{B}^{\prime}$, created by the above procedure up to the point where all cuts have been processed that appear in the list before the first $s,t$-cut $D^{*}_{st}$. By construction, $\mathcal{B}^{\prime}$ is linearly independent and each cutset $\partial U_{i}\in\mathcal{B}^{\prime}$ corresponds to a cut $(U_{i},\bar{U}_{i})$, where we may assume without loosing generality that $s,t\in U_{i}$. By the same argument as in the proof of Lemma 2.1, $\mathcal{B}^{\prime}$ cannot generate an $s,t$-cut as an $\oplus$-sum. Thus $\mathcal{B}^{\prime}\cup\{D^{*}_{st}\}$ is again linearly independent. In the next step, the canonical greedy algorithm therefore includes $D^{*}_{st}$ in the growing minimal weight basis $\mathcal{B}^{\prime}$. It follows that $D^{*}_{st}$ is relevant. $\triangleleft$
Claim: If $\partial V^{\prime}$ is not a minimum weight cut for any pair of vertices $s$ and $t$, then it is not relevant.
Let $T^{*}$ be an arbitrary Gomory-Hu tree for $G$. According to Eq. (2), $\partial V^{\prime}$ is a linear combination of cutsets $\partial V_{e}$ defined by $T^{*}$, where the edges $e$ separate vertices $s\in V_{e}$ and $t\in\bar{V}_{e}$. By (G2) and (G3), the weight $\lambda_{e}$ of the edge $e$ in $T^{*}$ satisfies $\lambda_{e}=|\partial V_{e}|$ and, by construction, is the minimum weight of cuts separating $s$ and $t$ if $\lambda_{e}$ is the minimal weight encountered along the unique path $P_{st}$ in $T^{*}$. Since $\partial V^{\prime}$ is not a minimum weight cut for any pair of vertices, it is in particular not a minimum weight cut for $s$ and $t$. Thus we have $|\partial V_{e}|<|\partial V^{\prime}|$ for every edge $e$ of the $T^{*}$ appearing in Eq. (2). Thus $D$ is a linear combination of strictly shorter cuts. By Lemma 3.2, therefore, $\partial V^{\prime}$ it is not relevant. $\triangleleft$
Taken together, a cut $D\in\mathfrak{D}$ is relevant if and only if it is a minimum weight cut for at least one pair of vertices $s,t\in V$.
There are graphs with an exponential number of minimal $s,t$-cuts. For instance the $K_{2,n-2}$ with unit edge-weights and $\{s,t\}$ for one part of the bipartition [22]. Since all these cuts are relevant by Thm. 3.6, the number of relevant cuts, like the number of relevant cycles [30], may also grow exponentially with the size of the graph $G$. In the unweighted case, i.e., $w_{pq}=1$ for all edges $pq\in E$, the number of relevant cuts is bounded by $|E|^{\Delta}$, where $\Delta$ is the maximum degree because every minimal $s,t$-cut contains at most $\Delta$ edges and there are no more than $O(|E|^{\Delta})$ distinct edge sets with at most $\Delta$ edges.
This argument does not generalize to arbitrary weights, however: Consider the graph $G$ obtained from two (potentially identical) instances of binary trees with $n$ leaves by identifying corresponding leaves. Thus $G$ has 3n23n-2 vertices and a maximum vertex degree of 33. Suppose all edge weights are unity except for edges incident with the leaves, which have weight $1/(2n)$. Denoting the roots of the trees by $s$ and $t$, every cut that contains an edge not incident with a tree-leaf has weight >1>1, while cutting exactly one of the two edges incident with a leaf yields an $s,t$-cut with weight $1/2$. There are $2^{n}$ distinct cuts of this type, all of which a minimum weight $s,t$-cut (see Fig. 1).
st1110.10.10.10.10.10.10.10.10.1110.11
Figure 1: Example of a graph with arbitrary weights and an exponential number of $s,t$-cuts. Following the construction in the main text joining two binary trees (solid vs dashed lines) with $n=5$ leaves, one edge adjacent per leave node has to be included in each minimal cut, yielding $2^{5}$ possibilities. Two such minimal cuts are marked in blue and red.
4 Algorithmic Considerations
As mentioned in the preliminaries, the set of all minimal $s,t$-cuts, which we have shown in Thm. 3.6 to coincide with the set of relevant cuts, can be computed from a Gomory-Hu tree $T^{*}$ and the directed acyclic graphs introduced by Picard and Queyranne [22] for the edges $uv$ of $T^{*}$. Both the Gomory-Hu tree $T^{*}$ (for recent advances, see [1, 2]) and the PQ-DAG $Q_{s,t}$ for any pair $s,t$ can be computed in polynomial time.
From this combined data structure, all the connectivity cuts, i.e., cuts with capacity equal to the connectivity value, can be listed with $O(|E||V|)$ effort [12]. This is feasible as there are at most $\binom{|V|}{2}$ such cuts. This is reminiscent of the polynomially many “prototypes” for relevant cycles described in [30], from which all relevant cycles can be listed. There is an important difference, however: For relevant cycles, there is a polynomial-time algorithm to count their number based on counting a constrained set of shortest paths [30]. In contrast, counting the number of minimal $s,t$-cuts is #P-complete [24], and thus at least as hard as NP-complete decision problems [28].
By Thm. 3.6, the most direct approach to computing the set of relevant cuts is to compute, for all distinct pairs $s,t\in V$, the set of all minimal $s,t$-cuts. We shall see, however, that the enumeration of all relevant cuts can be sped up considerably with the help of observations made by Gusfield and Noar in [12]. First we need some more notation.
For any $u,v\in V$, we say that $v$ is a successor of $u$ in $Q_{s,t}$, if either $u$ and $v$ belong to the same node in $Q_{s,t}$ or if the node in $Q_{s,t}$ that contains $v$ is a successor of the node in $Q_{s,t}$ that contains $u$. Let $Q_{s,t}(X,Y)$ with $X,Y\subseteq V$ and $X\cap Y=\emptyset$ denote the PQ-DAG resulting from $Q_{s,t}$ by contracting the set $X$ with all its successors, and contracting the set $Y$ with all its predecessors.
Lemma 4.1.
Any closed set in $Q_{s,t}(X,Y)$ that contains $s$ but not $t$ is a closed set in $Q_{s,t}$ that contains $\{s\cup X\}$ but not $\{t\cup Y\}$.
Proof 4.2.
By construction, $s$ is contracted into a super-node with all its successors and $t$ is contracted into a super-node with all its predecessors to create $Q_{s,t}$. Accordingly, the super-nodes containing $s$ and $t$ are the only vertices with out-degree 0 and in-degree 0, respectively. For any vertex $u\in V$, $s$ is a successor of $u$ and $t$ is a predecessors of $u$, as a consequence of this construction. Contracting $X$ with all its successors therefore includes $s$ but not $t$ as a successor. Analogously, contracting $Y$ with all its predecessors includes $t$ but not $s$ as a predecessor. Therefore, the super-node of out-degree 0 in $Q_{s,t}(X,Y)$ represents $\{s\cup X\}$, and the super-node of in-degree 0 represents $\{t\cup Y\}$. It follows that any closed set in $Q_{s,t}(X,Y)$ contains $\{s\cup X\}$ thereby $s$ but not $\{t\cup Y\}$ thereby $t$.
Lemma 4.3.
[12, Thm. 2.1] For a fixed pair of vertices $s,t\in V$, let $(u_{1},v_{1}),\ldots,(u_{k},v_{k})$ be the $k$ minimum-weight tree edges along the path between $s$ and $t$ in $T^{*}$. A cut $(V^{\prime},V^{\prime\prime})$ is a minimum $s,t$-cut if and only if it appears (as a closed set) in at least one PQ-DAG $Q_{u_{i},v_{i}}(s,t)$ for some $1\leq i\leq k$, where $Q_{u_{i},v_{i}}(s,t)$ is the graph resulting from contracting all successors of $s$ and all predecessors of $t$ in $Q_{u_{i},v_{i}}$, respectively, into a single node.
Theorem 4.4.
A cut $D\in\mathfrak{D}$ is relevant if and only if it is represented by at least one closed vertex set in a PQ-DAG $Q_{u,v}$ derived from an edge in the Gomory-Hu tree $T^{*}$, hence $\{u,v\}\in T^{*}$.
Proof 4.5.
$\Rightarrow$” By Thm. 3.6, if a cut is relevant, it is a weight-minimal $s,t$-cut for some pair of distinct vertices $s$ and $t$. By Lemma 4.3 any weight-minimal cut between two vertices $s$ and $t$ is represented by a closed set in exactly one PQ-DAG $Q_{u,v}(s,t)$, where $uv$ is a weight minimal edge along the path from $s$ to $t$ in $T^{*}$. By Lemma 4.1 any $u,v$-cut in $Q_{u,v}(s,t)$ is also a cut in $Q_{u,v}$.
$\Leftarrow$” Each edge $e=uv$ in a Gomory-Hu tree $T^{*}$ is weight minimal for at least the vertex pair $u$ and $v$ on the path consisting of only the edge $uv$ itself. The PQ-DAG $Q_{u,v}$ represents all weight minimal cuts between $u$ and $v$ as closed sets. The closed sets are therefore weight minimal cuts for two distinct vertices $s$ and $t$ and hence relevant as a consequence of Thm. 3.6.
Therefore, it is sufficient to compute only the PQ-DAGs associated with edges of a fixed Gomory-Hu tree $T^{*}$ of $G$. As the edges of a Gomory-Hu tree are labeled by the weights of associated minimal cuts, it is trivial to enumerate the relevant cuts in non-decreasing order using this method.
In addition to the approach of [12] and its derivation, several other algorithms for listing relevant cuts are feasible. By Lemma 3.4, one could use the bonds of $G$ as a candidate set and extract relevant cuts using the canonical greedy algorithm similar to Horton’s approach to computing minimum cycle bases [14]. Very recently, an algorithm for listing all $(s,t)$-bonds with near linear delay [25] has become available. A disadvantage is that the set of bonds must first be generated and then sorted by non-decreasing weight.
A more appealing alternative is to use one of the approaches for ordered generation of cuts. The first method of this type was proposed in [29]. Later an improved version based on a different solutions of maximum flow problems was published [32]. Since these algorithms produce cuts in the order of non-increasing weights, it is not difficult to determine whether a cut $(V^{\prime},\bar{V^{\prime}})$ is a minimum weight cut for any pair $s\in V$ and $t\in\bar{V}$. This can be done by keeping a matrix $M$, initialized by $M_{st}=\infty$ for all $s,t$, and updated by the following rule: if $M_{u,v}=\infty$, $u\in V$ and $v\in\bar{V^{\prime}}$ then $M_{u,v}\leftarrow|\partial V^{\prime}|$. In particular, because of Thm. 3.6, it is not necessary to check for linear independence of a cut with the set of all cuts of strictly smaller weight as is the case for minimal cycle bases. Moreover, the generation process can be terminated as soon as (i) minimum weight cuts have been observed for all vertices, i.e., upon generation of $(V,\bar{V}^{\prime})$, no entries $\infty$ were present in $M$, and (ii) the weight of the current cut is strictly larger than the weight of the previous one, which coincides with the largest finite entry in $M$. The enumeration algorithm in [32] has $\tilde{O}(|V||E|)$ delay, which dominates the $O(|V|^{2})$ effort for “filtering out” cuts that are not minimal $u,v$-cuts for any pair $u,v$.
5 Computational Experiments
Figure 2: Algorithmic performance for random graphs, partitioned by edge probability in the Erdős–Rényi model. Running times were tracked and related to the number of nodes (left) and the number of edges (right) of graphs. Lines represent the average running time, while environments show the 95% percentile interval. All plots present logarithmic y-scales. Regardless of edge density, methods show similar pattern.
Figure 3: Algorithmic performance for chemical graphs. Running times were tracked and related to the number of nodes (left) and the number of edges (right) of graphs. For improved resolution, chemical graphs containing 20 or less atoms are plotted independently (above), contrasting the full dataset (below). Lines represent the average running time, while environments show the 95% percentile interval. Please note that only the lower graphs follow a logarithmic y-scale.
Table 1: Summary of benchmarked chemical graphs. GUS-T performs best for the vast majority of cases, with exception of few very small graphs.
Category GUS-T GUS-P YEH
Total Running Time [s] 8340.50 35035.97 18616.66
Fastest Method [# Graphs] 9916 84 0
We implemented both methods derived from [12], enumerating closed sets in PQ-DAGs for all pairs of vertices (GUS-P) and only for PQ-DAGs for edges of $T^{*}$ (GUS-T), as well as the algorithm for enumerating cuts by non-increasing weights proposed by [32] (YEH). All methods where implemented in Python (v3.11) utilizing the ubiquitous graph library NetworkX (v. 3.3). All benchmarks where run on a consumer PC (Intel i9-9900K) with sufficient memory and without parallelization. Memory usage overall was negligible for all tested instances and is therefore not reported.
Benchmarks were performed on both artificial and realistic data. As generating cuts is of particular interest in the domain of computational chemistry, we used $10,000$ individual chemical graphs from the USPTO-10K dataset [21], a curated subset of reactions reported in the United States Patent and Trademark Office. Graphs ranged in size between 4 and 198 nodes with a median of 30 at 3 to 211 edges and a median of 33. As such, chemical graphs are considered to be sparse. Random graphs were constructed following the classic Erdős–Rényi model [10], generating 1000 valid instances per parameter variant, i.e. per each 6 to 105 nodes (at step size 2 until 40, then at step size 5) and per edge probability 0.25, 0.5, and 0.75. Only connected graphs where considered. The cut basis was computed for each method and dataset, logging individual running times. Times were averaged over $3$ iterations for chemical graphs.
Random graphs and chemical graphs both show a consistent, but markedly different hierarchy between methods. On all random graphs (see Fig. 2), GUS-T consistently outperformed both GUS-P and YEH, while GUS-P consistent outperformed YEH. Margins widen at increasing edge and node counts. Edge probability does not seems to have any influence on this relationship.
In chemical graphs, unsurprisingly, GUS-T also consistently outperformed both GUS-P and YEH for all but very small graphs (see Fig. 3, Tbl. 1). However, consistently, but more pronounced for larger graphs, YEH was significantly faster than GUS-P. Although GUS-P was the fastest algorithm for almost 1% of the graphs, manual inspection showed that this observation is most likely an artifact caused by fluctuations in loading times for very small graphs. Overall, YEH was a factor 2.2 slower then GUS-T to create a cut basis.
6 Concluding Remarks
With GUS-T, we derived an effective algorithm to compute exactly all relevant cuts of a graph, based on two central observations: relevant cuts coincide with the cuts that are weight-minimal for some pair of vertices, which in turn are fully encoded as the cuts implied by the edges of a Gomory Hu tree. Picard-Queyranne DAGs yield an effective framework towards this goal. GUS-T consistently outperformed other methods in our benchmarks, regardless of graph properties. Nevertheless, it should be acknowledged that YEH is overall more versatile when enumerating cuts.
GUS-T is in particular well suited to enumerate relevant cuts in chemical applications, where cuts relate to chemical reaction mechanism, such as for cut-and-join crossover operators [7]. It remains an open question, however, whether the restriction of cuts to the set of relevant cuts remains sufficient to ensure global reachability in the space of chemical graphs in conjunction with degree-preserving or even more restricted join operations.
Availability
A reference implementation of the methods developed in this paper are available via GitHub https://github.com/TGatter/Cut-Base-Toolkit.
Author Contributions
ND, TG, RG, and PFS designed the study, obtained the theoretical results and drafted the manuscript. ND and TG created the reference implementation and executed the benchmarks. The students of the Advanced Bioinformatics Course 2024 implemented [32] as a part of a course project.
Acknowledgements
This work was supported by the German Research Foundation (DFG) within SPP2363 (460865652) and the Novo Nordisk Foundation as part of MATOMIC (0066551). Research in the Stadler lab is supported by the German Federal Ministry of Education and Research BMBF through DAAD project 57616814 (SECAI, School of Embedded Composite AI). We gratefully acknowledge the contributions of the students of the Advanced Bioinformatics Course 2024: Marie Freudenberg, Julius Guntrum, Florian Kühn, Lukas Marche, Louisa Marie von Menges, and Kiet Ngo Tuan.